OpenStudy (anonymous):

Please help with this rational equation! 3/x^2+2/x=10/2x^2

OpenStudy (anonymous):

\[\frac{ 3 }{ x ^{2} }+\frac{ 2 }{ x } = \frac{ 10 }{ 2x ^{2} }\]

OpenStudy (isaiah.feynman):

What do you need to do!?

OpenStudy (anonymous):

I need to find x

OpenStudy (anonymous):

I'm really trying. The LCM is 2x^2 but I don't know how to multiply it by each side of the equation. That's the part where I get confused

OpenStudy (isaiah.feynman):